# Automated Machine Learning Pipeline for Training and Analysis Using Large Language Models

### v1

@@ 3.1 MLIPs for Acridine
We apply the new AMLP pipeline to acridine, a molecular crystal known to exhibit nine distinct crystal structures: eight polymorphs and one hydrate.[44] The primary challenge arises from their small lattice energy differences, typically within just a few kJ/mol.[45] Consequently, we seek to develop an MLIP capable of accurately distinguishing among the subtle structural variations of acridine polymorphs, thereby significantly improving predictive reliability compared to empirical force fields. All pure acridine polymorphs possess monoclinic symmetry, with experimental structures available from the Cambridge Structural Database (CSD) measured at room temperature. Due to variations in polymorph notation across different studies, we consistently adopt the naming convention provided by the CSD, as summarized in Table 1.
@@
Numerous theoretical studies have sought to determine the relative stability of acridine polymorphs.[44] However, classical force fields lack the flexibility required to accurately capture both the kinetic stability and the transitions between polymorphic forms.[46] Density Functional Theory (DFT) methods, while more accurate, still exhibit lattice-energy errors ranging from approximately 4.18 to 20 kJ/mol depending on the functional employed, whereas experimental energy differences between polymorphs rarely exceed 10 kJ/mol.[45] The MLIP developed here is trained on DFT data to reproduce the PES, thereby enabling simulations at larger system sizes and finite temperatures to explore polymorphic stability.
@@ 3.1.1 Dataset generation with LLMs
Based on these considerations, we performed DFT calculations using the Perdew-Burke-Ernzerhof (PBE) functional[47] and a plane-wave basis set, as implemented in VASP, both of which were recommended by AMLP. The agents also suggested an electronic convergence threshold of $10^{-6}\,\mathrm{eV}$, which we adopted. For the plane-wave cutoff, AMLP proposed 500 eV; here, we employed a more tighter value of 850 eV to ensure accurate total-energy evaluations. With respect to Brillouin-zone sampling, no explicit recommendation was provided by the agents, and we therefore selected a $\Gamma$-centered $7\times 7\times 7$ Monkhorst–Pack mesh.[48] For the treatment of long-range dispersion interactions, AMLP advised at least the use of D3(BJ) corrections[49]; however, we opted for Grimme’s D4 scheme,[50] which achieves accuracy comparable to the Many-Body Dispersion (MBD) method[51, 52, 53] while being substantially more efficient. The D4 correction ensures a reliable description of van der Waals interactions, which are essential for capturing the intermolecular forces governing crystal stability. All computational parameters are detailed in the SI. The crystal structures used in this study are presented in Table 1.
The DFT calculations of the eight polymorphs were subsequently processed using the AMLP code. The framework not only stores all relevant information in a structured .json format, as described in Section 2.1.3.
@@
Following the cell optimizations, we carried out AIMD simulations with VASP across a temperature range of 300–800 K. This temperature range was specified interactively by using AMLP, to capture both room-temperature and high-temperature dynamical behavior, thereby ensuring that the potential remains transferable even under conditions approaching or exceeding the melting temperature. framework. Specifically, the code identifies the relevant .json file from the preceding cell-optimization step, extracts the most stable geometry, and prepares the corresponding AIMD setup. Users can specify the temperature range and additional simulation parameters (e.g., timestep, ensemble), upon which AMLP automatically generates the corresponding input files. The analysis modules also provide default settings, ensuring that meaningful simulations can be carried out even without extensive user intervention.
For the present work, we employed the canonical (NVT) ensemble with a timestep of 1.0 fs over 10,000 steps, using a Langevin thermostat to regulate the temperature. In total, AMLP automatically generated 58 simulation directories, each containing the required VASP input files (INCAR, POSCAR, and KPOINTS) for 8 polymorphs evaluated at 7 distinct temperatures. These directories were fully prepared for direct submission to the computing environment, thereby eliminating the need for manual intervention during the setup stage. The only action required from the user is to provide the initial POTCAR file. From each AIMD trajectory, several configurations per polymorph were extracted, resulting in a comprehensive dataset giving both equilibrium and non-equilibrium structures as depicted in Figure 3.
Figure 3: Distribution of the different structures coming from the DFT with (a) the energy and (b) the maximum forces.
@@
We observe a bimodal distribution, with equilibrium configurations characterized by forces in the range of 0-0.5 eV/Å, and non-equilibrium configurations spanning approximately 2-10 eV/Å. Before training, we employed the AMLP code to convert the generated .json files into the appropriate dataset format (HDF5). The dataset was then automatically partitioned into training and validation subsets, using an 85 % to 15 % split by default, with a batch size of 4. Periodic boundary conditions (PBC) were identified and applied automatically during this process. To improve data quality, a force cutoff of 8 eV/Å was imposed to filter out outliers. After this preprocessing, a total of 8,108 structures remained available for model training.
@@ 3.1.2 Computational Details
Models were trained with the MACE implementation.[14] At present, users manually specify all training hyperparameters. In future releases, we will integrate LLM agents to assist with parameter selection and configuration generation. A cutoff radius of 6 Å was employed to define the local environment of each atom, ensuring the inclusion of relevant neighbors while maintaining computational efficiency.
The log loss function was chosen for its balance between sensitivity to small deviations and robustness against larger errors. The MACE fine-tuning with the foundation model mace-mpa-0-medium.model has been used. We used a foundation model, as it has been trained on the whole periodic table and on materials, making our model more flexible. Moreover, training from this checkpoint would speed up the convergence of our fitting process.[15]. We trained the models for a total of 350 epochs in two stages. During the first stage (250 epochs), equal weights were assigned to the energy and force terms. In the second stage (100 epochs), the weight on the forces was increased to be ten times larger than that on the energies, in order to prioritize force fitting. All details of the training parameters are provided in the SI.
@@ 3.1.4 Evaluation
When comparing the relative lattice energies of the acridine polymorphs, we observe a comparable trend between the committees and the reference DFT calculations. In the single-point evaluations, all three committees reproduce the general energetic ordering of the polymorphs, with some quantitative deviations depending on the model. Importantly, once the polymorphs are further relaxed with geometry optimization, the agreement with DFT improves systematically. The MLIPs seem to be more successful in reproducing the relative shape of the polymorphic energy landscape than in capturing absolute lattice energies.
@@
Figure 5: Comparison of relative lattice energies ($\Delta E_{\mathrm{lattice}}$) of acridine polymorphs computed with different committees (MACE A/B/C) against DFT fully relaxed structures. (a) shows single-point evaluations, while (b) corresponds to fully relaxed geometries.
To further assess structural fidelity, we examined the RMSD of atomic positions between DFT-relaxed structures and those obtained by geometry optimizations using the MACE models (Table 2). RMSDs were calculated after optimal structural alignment using the Kabsch algorithm, which determines the best-fit rotation and translation to superimpose the structures before computing deviations.[54, 55] All models achieve mean RMSDs of $\sim 0.048$$\text{\AA }$, indicating that the predicted and reference structures are essentially identical. MACE-C is slightly more consistent, showing lower standard deviation and maximum deviation values.
@@
Subsequently, we assessed the dynamical robustness of the trained models by performing microcanonical (NVE) simulations of 40 ps, including 1 ps equilibration, to evaluate energy conservation across the different polymorphs. The corresponding cumulative energies are summarized in Figure 6. For these simulations, all unit cells were replicated, reaching approximately 25 Å in each direction and resulting in system sizes ranging from 1,104 to 2,944 atoms, depending on the polymorph. To achieve this, a cell-replication tool was implemented within the AMLP-A module, enabling users to easily apply scaling factors along specified directions.
Figure 6: Cumulative energy conservation from NVE simulations of 39 ps following 1 ps of equilibration across all eight acridine polymorphs. Each panel corresponds to one polymorph, with results shown for three independently trained MLIPs: MACE-A (orange), MACE-B (green), and MACE-C (red). Solid lines represent the cumulative energy conservation, while shaded fluctuations reflect instantaneous deviations. All simulations show energy conservation on the order of $10^{-4}$.
In summary, these results confirm that the MLIPs trained within our AMLP framework provide reliable energy conservation in microcanonical simulations. With energy conservation in the $10^{-4}$ range, the committees establish a robust foundation for subsequent NVT simulations and analysis of radial distribution functions (RDFs), ensuring that observed structural and thermodynamic trends are not artifacts of poor energy conservation.
To investigate the structural stability of the different polymorphs, we performed NVT simulations using the different MLIP models. Within AMLP-A, users can select among various thermostats and integrators (see Section 2.1.6). At the end of each simulation, AMLP-a automatically computes RDFs for selected atomic pairs, thereby enabling a systematic analysis of the structural evolution with temperature. In particular, C–N pairs were analyzed as a probe of intramolecular stability, while N–N pairs served to characterize intermolecular packing interactions across a temperature range of T=300700T=300-700 K. The AMLP analysis tool was used to automatically generate the RDF plots shown in Figure 7, enabling direct comparison between different models.
@@
Figure 8: $P_{2}$ orientational order for acridine polymorphs at different temperatures using MACE-A and MACE-B potentials.
@@ 4 Conclusion
Our results show that the MLIPs automatically fine-tuned/trained with AMLP, achieves high accuracy, yielding close agreement with DFT references. Relative lattice energies is in agreement with DFT, rank-based stability trends are preserved. To further improve long-range accuracy, we plan to incorporate latent Ewald summation corrections into the AMLP pipeline, which explicitly address long-range electrostatics.[56]
@@
These findings emphasize the importance of dynamical validation, as equilibrium accuracy alone may be insufficient to guarantee model robustness at finite temperature. The AMLP pipeline simplifies this entire process, lowering the barrier of entry for researchers who are new to machine-learned interatomic potentials.
The current AMLP framework represents a first complete version including the most important functionalities for an automated fine-tuning and fitting of MLIPs. Future developments will focus on expanding usability, for instance, by enabling LLM agents to directly generate plots or analyze .json datasets on demand, as recently demonstrated for agentized scientific workflows.[57] We also plan to extend support beyond MACE to other state-of-the-art models such as NequIP[58], TorchMD[59], and FeNNol[60], thereby broadening AMLP’s applicability and robustness in computational materials science and theoretical chemistry.
@@ Acknowledgments
Supplementary Materials
The supplementary material is included below.
@@
Automated Machine Learning Pipeline for training and analysis using Large language models
Adam Lahouari,1 Jutta Rogal,1,2 and Mark E. Tuckerman1,3,4,5,6
1New York University, Department of Chemistry, New York, New York 10003, USA
2Initiative for Computational Catalysis, Flatiron Institute, New York, New York 10010, USA
3New York University, Department of Physics, New York, New York 10003, USA
4Courant Institute of Mathematical Sciences, New York University, New York, New York 10012, USA
5NYU-ECNU Center for Computational Chemistry at NYU Shanghai, Shanghai 200062, China
6Simons Center for Computational Physical Chemistry, Department of Chemistry, New York 10003, USA
E-mail: al9500@nyu.edu
@@
All DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP) with the Projector Augmented Wave (PAW) method.[17, 18, 19, 20] In this study on acridine, we decided to use VASP with the following parameters:

### v2

@@ 3.1 MLIPs for Acridine
We apply the new AMLP pipeline to acridine, a molecular crystal known to exhibit nine distinct crystal structures: eight polymorphs and one hydrate.[review_schur] The primary challenge arises from their small lattice energy differences, typically within just a few kJ/mol.[distribution_lattice_energy] Consequently, we seek to develop an MLIP capable of accurately distinguishing among the subtle structural variations of acridine polymorphs, thereby significantly improving predictive reliability compared to empirical force fields. All pure acridine polymorphs possess monoclinic symmetry, with experimental structures available from the Cambridge Structural Database (CSD) measured at room temperature. Due to variations in polymorph notation across different studies, we consistently adopt the naming convention provided by the CSD, as summarized in Table 1.
@@
Numerous theoretical studies have sought to determine the relative stability of acridine polymorphs.[review_schur] However, classical force fields lack the flexibility required to accurately capture both the kinetic stability and the transitions between polymorphic forms.[polymorphisme_md-qc] DFT, while more accurate, still exhibit lattice-energy errors ranging from approximately 4.18 to 20 kJ/mol depending on the functional employed, whereas experimental energy differences between polymorphs rarely exceed 10 kJ/mol.[distribution_lattice_energy] The MLIP developed here is trained on DFT data to reproduce the PES, thereby enabling simulations at larger system sizes and finite temperatures to explore polymorphic stability.
@@ 3.1.1 Dataset generation with LLMs
Based on these considerations, we performed DFT calculations using the Perdew-Burke-Ernzerhof (PBE) functional [PBE] and a plane-wave basis set, as implemented in VASP, both of which were recommended by AMLP. In this particular case, employing a plane-wave basis set is preferable over Gaussian basis sets when investigating crystalline structures due to its better suitability using PBC. Note that, while plane-wave implementations of hybrid density functionals have traditionally been computationally demanding compared to their use in GGA functionals such as PBE, recent methodological advances have significantly improved the efficiency of evaluating hybrid functionals for periodic systems in a plane-wave basis[mandal2022hybrid, mandal2019speeding, mandal2020efficient, mandal2018enhanced], making the use of hybrid functionals more feasible for running simulations. The agents also suggested an electronic convergence threshold of $10^{-6}\,\mathrm{eV}$, which we adopted. For the plane-wave cutoff, AMLP proposed 500 eV; here, we employed a more tighter value of 850 eV to ensure accurate total-energy evaluations. With respect to Brillouin-zone sampling, no explicit recommendation was provided by the agents, and we therefore selected a $\Gamma$-centered $7\times 7\times 7$ Monkhorst–Pack mesh.[monkhorst] For the treatment of long-range dispersion interactions, AMLP advised at least the use of D3(BJ) corrections[d3bj]; however, we opted for Grimme’s D4 scheme,[grimm-d4] which achieves accuracy comparable to the Many-Body Dispersion (MBD) method[mbd-tatch1, mbd-tatch2, mbd-tatch3] while being substantially more efficient. The D4 correction ensures a reliable description of van der Waals interactions, which are essential for capturing the intermolecular forces governing crystal stability. All computational parameters are detailed in the SI. The crystal structures used in this study are presented in Table 1.
The DFT calculations of the eight polymorphs were subsequently processed using the AMLP code. The framework stores all relevant information in a structured .json format, as described in step 3 of Section 2.2.
@@
Following the cell optimizations, we carried out AIMD simulations with VASP across a temperature range of 300–800 K. This temperature range was specified interactively by using AMLP, to capture both room-temperature and high-temperature dynamical behavior, thereby ensuring that the potential remains transferable even under conditions approaching or exceeding the melting temperature. Specifically, the code identifies the relevant .json file from the preceding cell-optimization step, extracts the most stable geometry, and prepares the corresponding AIMD setup. Users can specify the temperature range and additional simulation parameters (e.g., timestep, ensemble), upon which AMLP automatically generates the corresponding input files. The analysis modules also provide default settings, ensuring that meaningful simulations can be carried out even without extensive user intervention.
For the present work, we employed the canonical (NVT) ensemble with a timestep of 1.0 fs over 10,000 steps, using a Langevin thermostat to regulate the temperature. These parameters are user-defined and can be adjusted based on the specific requirements of each system. AMLP provides sensible default settings to serve as initial starting points for users. Additionally, users can consult the integrated LLM agents for system-specific parameter recommendations tailored to their particular molecular system. In total, AMLP automatically generated 58 simulation directories, each containing the required VASP input files (INCAR, POSCAR, and KPOINTS) for 8 polymorphs evaluated at 7 distinct temperatures. These directories were fully prepared for direct submission to the computing environment, thereby eliminating the need for manual intervention during the setup stage. The only action required from the user is to provide the initial POTCAR file. From each AIMD trajectory, several configurations per polymorph were extracted, resulting in a comprehensive dataset giving both equilibrium and non-equilibrium structures as depicted in Figure 3.
Figure 3: Distribution of the different structures coming from the DFT with (a) the energy per atom and (b) the maximum forces.
@@
We observe a bimodal distribution, with equilibrium configurations characterized by forces in the range of 0-0.5 eV/Å, and non-equilibrium configurations spanning approximately 2-10 eV/Å. Before training, we employed the AMLP code to convert the generated .json files into the appropriate dataset format (HDF5). The dataset was then automatically partitioned into training and validation subsets, using an 85 % to 15 % split by default, with a batch size of 4. PBC were identified and applied automatically during this process. To improve data quality, a force cutoff of 8 eV/Å was imposed to filter out outliers. After this preprocessing, a total of 8,108 structures remained available for model training.
@@ 3.1.2 Computational Details
Models were trained with the MACE implementation.[MACE] At present, users manually specify all training hyperparameters. In future releases, we will integrate LLM agents to assist with parameter selection and configuration generation. A cutoff radius of 6 Å was employed to define the local environment of each atom, ensuring the inclusion of relevant neighbors while maintaining computational efficiency.
The log loss function was chosen for its balance between sensitivity to small deviations and robustness against larger errors. The MACE fine-tuning with the foundation model mace-mpa-0-medium.model has been used. We used a foundation model, as it has been trained on the whole periodic table and on materials, making our model more flexible. Moreover, training from this checkpoint would speed up the convergence of our fitting process.[fine-tuning-juttafriend]. We trained the models for a total of 350 epochs in two stages. During the first stage (250 epochs), equal weights were assigned to the energy and force terms. In the second stage (100 epochs), the weight on the forces was increased to be ten times larger than that on the energies, in order to prioritize force fitting. For both stages of the training procedure, the same dataset was used. All details of the training parameters are provided in the SI.
@@ 3.1.4 Evaluation
When comparing the relative lattice energies of the acridine polymorphs, all three committees reproduce the DFT energetic ordering reasonably well, with model-dependent quantitative offsets. By contrast, the untuned foundation model (purple curve in Fig. 5) deviates markedly; for example, it places ACRDIN04 at  30 kJ/mol above the most stable form, which is larger than typical polymorphic separations. In single-point evaluations, the committees already recover the general trend, and subsequent geometry optimization further improves agreement with DFT for every model—including the MPA model—though the MPA model still exhibits larger relative-energy errors than the fine-tuned committees. Overall, these fine-tuned MLIPs better capture the shape of the polymorphic energy landscape (rank ordering and relative trends) than the absolute lattice energies.
@@
Figure 5: Relative lattice energies ($\Delta E_{\mathrm{lattice}}$) of acridine polymorphs predicted by the MPA–MACE foundation model and by fine-tuned committee models (MACE-A/B/C), compared against fully relaxed DFT references. (a): single-point evaluations on fixed geometries; (b): energies after geometry optimization.
To further assess structural fidelity, we examined the RMSD of atomic positions between DFT-relaxed structures and those obtained by geometry optimizations using the MACE models (Table 2). RMSDs were calculated after optimal structural alignment using the Kabsch algorithm, which determines the best-fit rotation and translation to superimpose the structures before computing deviations.[kabsh1, kabsh2] All models achieve mean RMSDs of $\sim 0.048$$\text{\AA }$, indicating that the predicted and reference structures are essentially identical. MACE-C is slightly more consistent, showing lower standard deviation and maximum deviation values.
@@
Subsequently, to assess the dynamical robustness of the trained models, we performed 40 ps microcanonical (NVE) simulations with 1 ps of equilibration to evaluate energy conservation across different polymorphs. All models demonstrated excellent energy conservation in the range of $10^{-4}$ range for all polymorphs (Figure S1). Further details on the evaluation methodology and complete results are provided in the Supporting Information.
These results confirm that the MLIPs trained within our AMLP framework provide reliable energy conservation in microcanonical simulations. The committees establish a robust foundation for subsequent NVT simulations and analysis of radial distribution functions (RDFs), ensuring that observed structural and thermodynamic trends are not artifacts of poor energy conservation.
To investigate the structural stability of the different polymorphs, we performed NVT simulations using the different MLIP models. Within AMLP-A, users can select among various thermostats and integrators (see Section 2.3). At the end of each simulation, AMLP-a automatically computes RDFs for selected atomic pairs, thereby enabling a systematic analysis of the structural evolution with temperature. In particular, C–N pairs were analyzed as a probe of intramolecular stability, while N–N pairs served to characterize intermolecular packing interactions across a temperature range of T=300700T=300-700 K. The AMLP analysis tool was used to automatically generate the RDF plots shown in Figure 6, enabling direct comparison between different models.
@@
Figure 7: $P_{2}$ orientational order for acridine polymorphs at different temperatures using (a) MACE-A and (b) MACE-B potentials.
@@
To assess the transferability of our MLIP to polymorphs outside the training set, we tested Form VIII, a recently characterized experimental structure that was not included in our dataset.[review_schur] Geometry optimization and MD simulations were performed in both NVE and NVT ensembles. The optimized structures yielded relative lattice energies within an acceptable range:  10.3 kJ/mol for MACE-A and MACE-B, and higher for MACE-C around  17 kJ/mol (see Figure S12a).The models also produced stable trajectories with proper energy conservation under the NVE ensemble (see Figure S12b) and stable dynamics under the NVT ensemble at 300 K for MACE-A and MACE-B (see Figure S12c). However, simulations at higher temperatures exhibited instabilities, indicating that the PES requires refinement for this polymorph under elevated-temperature conditions. This case study illustrates both the strengths and current limitations of the trained MLIPs. While the models shows reasonable transferability to unknown forms at low to moderate temperatures, configurations far from the training distribution—such as those accessed at high temperatures—require additional reference data. Crucially, AMLP facilitates this refinement process: structures from the unstable MD trajectories can be easily extracted, and the pipeline helps the users to generate the necessary single-point calculations or AIMD simulations to augment the training set. This demonstrates the practical value of the automated framework for iterative MLIPs development and extension to new chemical environments.
@@ 4 Conclusion
Our results show that the MLIPs automatically fine-tuned/trained with AMLP, achieves high accuracy, yielding close agreement with DFT references. Relative lattice energies is in agreement with DFT, rank-based stability trends are preserved. To further improve long-range accuracy, we plan to incorporate latent Ewald summation corrections into AMLP, which explicitly address long-range electrostatics.[long-rangecorrection-ML]
@@
These findings emphasize the importance of dynamical validation, as equilibrium accuracy alone may be insufficient to guarantee model robustness at finite temperature. AMLP simplifies this entire process, lowering the barrier of entry for researchers who are new to machine-learned interatomic potentials.
The current AMLP framework represents a first complete version including the most important functionalities for an automated fine-tuning and fitting of MLIPs. Future developments will focus on expanding usability, for instance, by enabling LLM agents to directly generate plots or analyze .json datasets on demand, as recently demonstrated for agentized scientific workflows.[DREAMSDensityFunctional2025] We also plan to extend support beyond MACE to other state-of-the-art models such as NequIP[nequip], TorchMD[TorchMD], and FeNNol[fennol], thereby broadening AMLP’s applicability and robustness in computational materials science and theoretical chemistry. In addition, the forthcoming release will introduce an automated AL orchestrator that translates user-specified objectives into cycling through inexpensive configuration generation with the use of different foundation models, selective high-level labeling, model retraining, and validation.
@@ Acknowledgments
Supporting Information
Multi-agent system outputs including key findings and computational recommendations; VASP parameters for cell optimization; AIMD control parameters and electronic structure settings; MACE machine learning potential training parameters and atomic reference energies; energy conservation analysis demonstrating simulation stability; radial distribution functions characterizing local structure for all atomic pairs; out-of-sample polymorph validation results.
TOC Graphic
@@
Automated Machine Learning Pipeline: Large Language Models-Assisted Automated Dataset Generation for Training Machine-Learned Interatomic Potentials.
Adam Lahouari1,∗, Jutta Rogal1,2, and Mark E. Tuckerman1,3,4,5,6
1NYU, Department of Chemistry, New York, NY 10003, USA
2Initiative for Computational Catalysis, Flatiron Institute, NY 10010, USA
3NYU, Department of Physics, New York, NY 10003, USA
4Courant Institute of Mathematical Sciences, NYU, NY 10012, USA
5NYU-ECNU Center for Computational Chemistry, Shanghai 200062, China
6Simons Center for Computational Physical Chemistry, NYU, NY 10003, USA
Correspondence: al9500@nyu.edu
@@
All DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP) with the Projector Augmented Wave (PAW) method.[vasp1, vasp2, vasp3, vasp4] In this study on acridine, we decided to use VASP with the following parameters:
@@
Energy Conservation
The corresponding cumulative energies are summarized in Figure S1. For these simulations, all unit cells were replicated, reaching approximately 25 Å in each direction and resulting in system sizes ranging from 1,104 to 2,944 atoms, depending on the polymorph. To achieve this, a cell-replication tool was implemented within the AMLP-A module, enabling users to easily apply scaling factors along specified directions.
Figure S1: Cumulative energy conservation from NVE simulations of 39 ps following 1 ps of equilibration across all eight acridine polymorphs. Each panel corresponds to one polymorph, with results shown for three independently trained MLIPs: MACE-A (orange), MACE-B (green), and MACE-C (red). Solid lines represent the cumulative energy conservation, while shaded fluctuations reflect instantaneous deviations. All simulations show energy conservation on the order of $10^{-4}$.
@@
Out-of-Sample Validation: ACRIDIN VIII
Figure S12: Performance evaluation of MACE committee models on ACRIDINVIII, a polymorph excluded from the training set. (a) Relative lattice energies of acridine polymorphs after geometry optimization, comparing DFT reference calculations with MPA-MACE and individual committee models (MACE-A, MACE-B, MACE-C). ACRIDINVIII (Form VIII) was not included in the training data. (b) Energy drift assessment during 40 ps NVE molecular dynamics simulations at 300 K using the three MACE committee models, demonstrating energy conservation with MACE-C showing the lowest drift ($\Delta E~2.0\times 10^{-6}$ eV/atom). (c) Temperature stability during 40 ps NVT simulations at 300 K for MACE-A and MACE-B committees, showing equilibration around the target temperature with minimal fluctuations.